One inch is equal to approximately 2.5 centimeters. Let x represent inches and y represent centimeters. Write an equation in standard form relating x and y. Give the values of A, B, and C.
Answer by mananth(16949)

2y=x
2y-x=0
-x+2y=0
...
A = -1
B=2
C=0